# Predictability Enables Parallelization of Nonlinear State Space Models

### v1

@@ Abstract
The rise of parallel computing hardware has made it increasingly important to understand which nonlinear state space models can be efficiently parallelized. Recent advances like DEER (Lim et al., 2024) or DeepPCR (Danieli et al., 2023) have shown that evaluating a state space model can be recast as solving a parallelizable optimization problem, and sometimes this approach can yield dramatic speed-ups in evaluation time. However, the factors that govern the difficulty of these optimization problems remain unclear, limiting the larger adoption of the technique. In this work, we establish a precise relationship between the dynamics of a nonlinear system and the conditioning of its corresponding optimization formulation. We show that the predictability of a system, defined as the degree to which small perturbations in state influence future behavior, impacts the number of optimization steps required for evaluation. In predictable systems, the state trajectory can be computed in $\mathcal{O}((\log T)^{2})$ time, where $T$ is the sequence length, a major improvement over the conventional sequential approach. In contrast, chaotic or unpredictable systems exhibit poor conditioning, with the consequence that parallel evaluation converges too slowly to be useful. Importantly, our theoretical analysis demonstrates that for predictable systems, the optimization problem is always well-conditioned, whereas for unpredictable systems, the conditioning degrades exponentially as a function of the sequence length. We validate our claims through extensive experiments, providing practical guidance on when nonlinear dynamical systems can be efficiently parallelized, and highlighting predictability as a key design principle for parallelizable models.
Equal contribution.
1 Department of Statistics, Stanford University, Stanford, CA, USA
2 Wu Tsai Neurosciences Institute, Stanford University, Stanford, CA, USA
3 Thomas J. Watson Research Center, IBM Research, Yorktown Heights, NY, USA
4 Almaden Research Center, IBM Research, San Jose, CA, USA
Correspondence should be addressed to X.G. (xavier18@stanford.edu), L.K. (leokoz8@gmail.com), and S.W.L. (scott.linderman@stanford.edu).
@@ 1 Introduction
Parallelization has been central to recent breakthroughs in deep learning, with GPUs enabling the fast training of large neural networks. In contrast, nonlinear state space models like recurrent neural networks (RNNs) often resist efficient parallelization on GPUs due to their inherently sequential structure.
Recent work addresses this mismatch by reformulating sequential dynamics into parallelizable optimization problems. Notably, the DeepPCR/DEER algorithm (Danieli et al., 2023; Lim et al., 2024) evaluates nonlinear state space dynamics by minimizing a residual-based merit function, facilitating efficient parallel computation via the Gauss-Newton method. [^1] DEER (Lim et al., 2024) and DeepPCR (Danieli et al., 2023) were concurrent works that both proposed to use the Gauss-Newton method for optimizing nonlinear sum of squares to parallelize sequential processes. In this paper, we therefore use DEER, DeepPCR, and Gauss-Newton interchangeably. Gonzalez et al. (2024) further developed these methods, including quasi-Newton methods and trust regions methods for parallel evaluation of nonlinear dynamical systems. These methods evaluate nonlinear dynamical systems by iteratively linearizing the nonlinear system and evaluating the resulting linear dynamical system (LDS) with a parallel (a.k.a. associative) scan (Stone, 1973; Blelloch, 1990). Each parallel evaluation of an LDS implements one optimization step  (Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024).
@@
Drawing from nonlinear dynamical systems theory—particularly contraction analysis (Lohmiller and Slotine, 1998) and Lyapunov exponent methods (Pikovsky and Politi, 2016)—we formalize the relationship between system predictability and the conditioning of the merit function. Unpredictable systems are dynamical systems whose future behavior is highly sensitive to small perturbations. A common example is a chaotic system, like the weather: a butterfly flapping its wings in Tokyo today can lead to a thunderstorm in Manhattan next month (Lighthill, 1986; Strogatz, 2018). By contrast, predictable systems are those in which small perturbations are “forgotten.” A familiar example is aviation: a patch of choppy air rarely makes an airplane land at the wrong airport. A more formal definition of (un)predictability is given in Definition 1. Our results establish key theoretical principles, make connections between optimization theory and dynamical systems, and demonstrate the practical applicability of parallel computations across a wide range of nonlinear state space modeling tasks.
@@
The paper is organized as follows. Section 2 provides background, with formal definitions of predictable and unpredictable nonlinear state space models. Section 3 presents two key theoretical results that characterize the conditioning of the merit function, showing that the Polyak–Łojasiewicz (PL) constant $\mu$ of the merit function is controlled by the predictability of the dynamics (Theorem 2), and that the Lipschitz constant of the residual function Jacobian is governed by the nonlinearity of the dynamics (Theorem 3). Section 4 then uses the results about the conditioning of the merit function to prove results about Gauss-Newton in particular. We prove global linear rates of convergence for Gauss-Newton, with the precise rate scaling with the unpredictability of the problem (Theorem 4), and we characterize the basin of quadratic convergence in terms of the predictability and nonlinearity of the underlying dynamics (Theorem˜5). In Section 5 we illustrate our results with experiments, and in Section 6 we conclude by summarizing context, implications, limitations, and future directions.
@@ 2 Problem Statement & Background
[…]
@@
Systems of the form (1) are widespread across essentially all fields of science and engineering. Examples include physics (numerical weather prediction, molecular dynamics), biology (gene regulatory networks, population dynamics), engineering (control, robotics), and economics (macroeconomic forecasting, asset pricing). In machine learning, sequential operations arise in recurrent neural networks, iterative optimization, and the sampling pass of a diffusion model (Song et al., 2021; Danieli et al., 2023; Tang et al., 2024). Sequential operations even appear in the problem of evaluating transformer blocks over depth (Dehghani et al., 2019; Schöne et al., 2025; Geiping et al., 2025; Calvo-González et al., 2025; ARC Prize Team, 2025). In probabilistic modeling, sequential operations arise in Expectation-Maximization and Markov Chain Monte Carlo (Zoltowski et al., 2025). In all of these cases, the state evolves through nonlinear transformations that capture the system’s underlying dynamics.
The obvious approach is to sequentially compute the states according to eq.˜1, taking $T$ steps. Alternatively, one can cast state evaluation as an optimization problem. While less intuitive, an advantage of this approach is that it admits parallel computation (Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024). Depending on the properties of the nonlinear state space model, the optimization algorithm, and the available hardware, the latter approach can be significantly faster than sequential evaluation.
[…]
@@
DeepPCR (Danieli et al., 2023) and DEER (Lim et al., 2024) minimize the merit function using Gauss–Newton updates. Each update takes the form
@@
Due to this block bidiagonal structure, solving $\mathbf{J}(\mathbf{s}^{(i)})^{-1}\,\mathbf{r}(\mathbf{s}^{(i)})$ amounts to solving a linear recursion, which can be done in $\mathcal{O}(\log T)$ time with a parallel scan (Blelloch, 1990; Martin and Cundy, 2018; Smith et al., 2023; Lim et al., 2024; Gonzalez et al., 2024). Further details are given in Appendix˜A.
@@
Predictability is usually defined through its antonym: unpredictability (Lighthill, 1986; Strogatz, 2018). In an unpredictable system, the system’s intrinsic sensitivity amplifies small perturbations and leads to massive divergence of trajectories. Predictable systems show the opposite behavior: small perturbations are diminished over time, rather than amplified. The notion of (un)predictability can be formalized through various routes such as chaos theory (Gleick, 2008; Schuster and Just, 2006) and contraction analysis (Lohmiller and Slotine, 1998; Bullo, 2024).
The definition of predictability comes from the Largest Lyapunov Exponent (LLE) (Pikovsky and Politi, 2016; Strogatz, 2018):
@@ 3.1 The Merit Function is PL
To begin, we show that the merit function (2) satisfies the Polyak-Łojasiewicz (PL) condition (Karimi et al., 2016), also known as the gradient dominance condition (Fazel et al., 2018). A function $\mathcal{L}(\mathbf{s})$ is $\mu$-PL if it satisfies, for $\mu>0$,
@@
See Appendix B. This result, known in the literature for general sum-of-squares (Nesterov and Polyak, 2006), is included here for context and completeness. ∎
Proposition˜1 is important as it characterizes the flatness of the merit function. If $\mu$ is very small in a certain region, this indicates that the norm of the gradient can be very small in that region, which can make gradient-based optimization inefficient. Proposition˜1 also links $\sigma_{\mathrm{min}}(\mathbf{J})$—important for characterizing the conditioning of $\mathbf{J}$—to the geometry of the merit function landscape.
@@ 3.2 Merit Function PL Constant is Controlled by the Largest Lyapunov Exponent of Model
The proof mechanism we have sketched upper and lower bounds $\|\mathbf{J}^{-1}\|_{2}$ in terms of norms of Jacobian products. We only use the assumption in eq.˜10 to express those bounds in terms of $\lambda$. As we discuss at length in Appendix˜C, we can use different assumptions from eq.˜10 to get similar results. Theorem˜2 and its proof should be thought of as a framework, where different assumptions (which may be more or less relevant in different settings) can be plugged in to yield specific results.
@@ 3.3 Residual function Jacobian Inherits the Lipschitzness of the Nonlinear State Space Model
[…]
@@ 4 Rates of Convergence for Optimizing the Merit Function
Theorem 4 also lets us establish key worst-case and average-case bounds on the number of steps needed for Gauss-Newton to converge to within a given distance of the solution. In particular, when $\chi_{w}$ does not depend on the sequence length $T$, then Theorem˜4 implies Gauss-Newton will only require $\mathcal{O}\left((\log T)^{2}\right)$ total computational time, with one $\log$ factor coming from the parallel scan at each optimization step and the other coming from the total number of optimization steps needed. We elaborate on these points in Appendix G.
@@
See Appendix H. We make no claim about the originality of lower bounding the size of the basin of quadratic convergence in Gauss-Newton. In fact, our proof of Theorem˜5 closely follows the convergence analysis of Newton’s method in Section 9.5.3 of Boyd and Vandenberghe (2004). Our contribution is we highlight the elegant way the predictability $\lambda$ and non-linearity $L$ of a dynamical system influence an important feature of its merit function’s landscape. ∎
In the next section, we show empirical results that support these theoretical findings.
@@ 5 Experiments
In this section we conduct experiments to support the theory developed above, demonstrating that predictability enables parallelization of nonlinear dynamical systems. To illustrate this point, we use Gauss-Newton optimization (DeepPCR (Danieli et al., 2023) aka DEER (Lim et al., 2024)). We provide more experimental details in Appendix J.
@@
Theorem 2 predicts a sharp phase transition in the conditioning of the merit function at $\lambda=0$, which should be reflected in the number of optimization steps required for convergence. To empirically validate this prediction, we vary both the LLE and sequence length $T$ within a parametric family of recurrent neural networks (RNNs), and measure the number of steps DEER takes to converge. We generate mean-field RNNs following Engelken et al. (2023), scaling standard normal weight matrices by a single parameter that controls their variance and therefore the expected LLE. In Figure˜2, we observe a striking correspondence between the conditioning of the optimization problem (represented by $\log\tilde{\mu}$, where $\tilde{\mu}$ is the lower bound for $\mu$ from Theorem˜2) and the number of steps DEER takes to convergence. This relationship holds across the range of LLEs, $\lambda$, and sequence lengths, $T$. There is a rapid threshold phenomenon around $\lambda=0$, which divides predictable from unpredictable dynamics, precisely as expected from Theorem˜2. As we discuss in Appendix J.1, the correspondence between $\log\tilde{\mu}$ and the number of optimization steps needed for convergence can be explained by DEER iterates approaching the basin of quadratic convergence with linear rate.
@@
Figure 2: Threshold phenomenon in DEER convergence based on system predictability. In a family of RNNs, DEER has fast convergence for predictable systems and prohibitively slow convergence for chaotic systems. Left (Theory): We depict Theorem 2, illustrating how the conditioning of the optimization problem degrades as $T$ and the LLE ($\lambda$) increase. Center (Experiment): We vary $\lambda$ across the family of RNNs, and observe a striking concordance in the number of DEER optimization steps empirically needed for convergence with our theoretical characterization of the conditioning of the optimization problem. Right: For 20 seeds, each with 50 different values of $\lambda$, we plot the relationship between $\lambda$ and the number of DEER steps needed for convergence for the sequence length $T=954$ (gray line in left and center panels). We observe a sharp increase in the number of optimization steps at precisely the transition between predictability and unpredictability.
@@
DEER may still converge quickly even if the system is unpredictable in certain regions. As long as the system is predictable on average, as indicated by a negative LLE, DEER can still converge quickly. This phenomenon is why we framed Theorem 2 in terms of the LLE $\lambda$ and burn-in constants $a$, as opposed to a weaker result that assumes the system Jacobians have singular values less than one over the entire state space (see our discussion of condition (10) vs. condition (21) in Appendix C).
@@
Notably, prior works such as Lim et al. (2024) and Gonzalez et al. (2024) initialized optimization from $\mathbf{s}^{(0)}=\mathbf{0}$, which lies entirely in the unstable region. Thus, our theoretical insights into predictability and parallelizability suggest practical improvements for initialization.
Figure 3: DEER converges quickly for Langevin dynamics in a two-well potential. (Left) An illustration of the two-well potential state space in $D=2$. We superimpose a contour plot of the potential on a color scheme showing the spectral norm of the dynamics Jacobian (blue indicates stability, red instability). (Center) A trace plot for the $y$-coordinate. The LLE of the system is 0.015-0.015. (Right) We observe that this system, which has negative LLE, enjoys sublinear scaling in the sequence length $T$ in the number of DEER iterations needed to converge. We plot the median number of DEER steps to convergence over 20 random seeds.
@@
Finally, we demonstrate a practical application of our theory in the efficient parallelization of chaotic observers. Observers are commonly used to reconstruct the full state of a system from partial measurements (Luenberger, 1979; Simon, 2006). On nine chaotic flows from the dysts benchmark dataset (Gilpin, 2021a), Table 1 shows that while DEER converges prohibitively slowly on chaotic systems, it converges rapidly on stable observers of these systems, in accordance with our theory that predictability implies parallelizability. For more details, see Appendix J.5.

### v2

@@ Abstract
The rise of parallel computing hardware has made it increasingly important to understand which nonlinear state space models can be efficiently parallelized. Recent advances like DEER [1] or DeepPCR [2] have shown that evaluating a state space model can be recast as solving a parallelizable optimization problem, and sometimes this approach can yield dramatic speed-ups in evaluation time. However, the factors that govern the difficulty of these optimization problems remain unclear, limiting the larger adoption of the technique. In this work, we establish a precise relationship between the dynamics of a nonlinear system and the conditioning of its corresponding optimization formulation. We show that the predictability of a system, defined as the degree to which small perturbations in state influence future behavior, impacts the number of optimization steps required for evaluation. In predictable systems, the state trajectory can be computed in $\mathcal{O}((\log T)^{2})$ time, where $T$ is the sequence length, a major improvement over the conventional sequential approach. In contrast, chaotic or unpredictable systems exhibit poor conditioning, with the consequence that parallel evaluation converges too slowly to be useful. Importantly, our theoretical analysis demonstrates that for predictable systems, the optimization problem is always well-conditioned, whereas for unpredictable systems, the conditioning degrades exponentially as a function of the sequence length. We validate our claims through extensive experiments, providing practical guidance on when nonlinear dynamical systems can be efficiently parallelized, and highlighting predictability as a key design principle for parallelizable models.
@@ 1 Introduction
Parallelization has been central to breakthroughs in deep learning, with GPUs enabling the fast training of large neural networks. In contrast, nonlinear state space models like recurrent neural networks (RNNs) have resisted efficient parallelization on GPUs due to their sequential nature.
Recent work addresses this mismatch by reformulating sequential dynamics into parallelizable optimization problems. Notably, the DEER/DeepPCR algorithm [1, 2] evaluates nonlinear state space dynamics by minimizing a residual-based merit function, facilitating efficient parallel computation
@@
via the Gauss-Newton method. [^1] DEER [1] and DeepPCR [2] were concurrent works that both proposed to use the Gauss-Newton method for optimizing nonlinear sum of squares to parallelize sequential processes. In this paper, we therefore use DEER, DeepPCR, and Gauss-Newton interchangeably. Gonzalez et al. [3] further developed these methods, including quasi-Newton methods and trust regions methods for parallel evaluation of nonlinear dynamical systems. These methods evaluate nonlinear dynamical systems by iteratively linearizing the nonlinear system and evaluating the resulting linear dynamical system (LDS) with a parallel (a.k.a. associative) scan [4, 5]. Each parallel evaluation of an LDS implements one optimization step  [1, 2, 3, 6].
@@
Drawing from nonlinear dynamical systems theory—particularly contraction analysis [7] and Lyapunov exponent methods [8]—we formalize the relationship between system predictability and the conditioning of the merit function. Unpredictable systems are dynamical systems whose future behavior is highly sensitive to small perturbations. A common example is a chaotic system, like the weather: a butterfly flapping its wings in Tokyo today can lead to a thunderstorm in Manhattan next month [9, 10]. By contrast, predictable systems are those in which small perturbations are “forgotten.” A familiar example is aviation: a patch of choppy air rarely makes an airplane land at the wrong airport. A more formal definition of (un)predictability is given in Definition 1. Our results establish key theoretical principles, make connections between optimization theory and dynamical systems, and demonstrate the practical applicability of parallel computations across a wide range of nonlinear state space modeling tasks.
@@
The paper is organized as follows. Section 2 provides background, with formal definitions of predictable and unpredictable nonlinear state space models. Section 3 presents two key theoretical results that characterize the conditioning of the merit function, showing that the Polyak–Łojasiewicz (PL) constant $\mu$ of the merit function is controlled by the predictability of the dynamics (Theorem 2), and that the Lipschitz constant of the residual function Jacobian is governed by the nonlinearity of the dynamics (Theorem 3). Section 4 then uses the results about the conditioning of the merit function to prove results about Gauss-Newton in particular. We prove global linear rates of convergence for Gauss-Newton, with the precise rate scaling with the unpredictability of the problem (Theorem 4), and we characterize the basin of quadratic convergence in terms of the predictability and nonlinearity of the underlying dynamics (˜5). In Section 5 we illustrate our results with experiments, and in Section 6 we conclude by summarizing context, implications, limitations, and future directions.
@@ 2 Problem Statement & Background
Throughout the paper, we use $T$ to denote the length of a sequence and $D$ to represent the dimensionality of a nonlinear state space model. Elements in $\mathbb{R}$, $\mathbb{R}^{D}$ or $\mathbb{R}^{D\times D}$ are written using non-bold symbols, while elements in $\mathbb{R}^{TD}$ or $\mathbb{R}^{TD\times TD}$ are denoted with bold symbols.
@@
Systems of the form (1) are widespread across science and engineering. Examples include physics (numerical weather prediction, molecular dynamics), biology (gene regulatory networks, population dynamics), engineering (control, robotics), and economics (macroeconomic forecasting, asset pricing). In machine learning, sequential operations arise in recurrent neural networks, iterative optimization, and the sampling pass of a diffusion model [2, 11]. Sequential operations even appear in the problem of evaluating transformer blocks over depth [12, 13, 14, 15, 16, 17]. In probabilistic modeling, sequential operations arise in Markov Chain Monte Carlo [18]. In all of these cases, the state evolves through nonlinear transformations that capture the system’s underlying dynamics.
The obvious approach is to sequentially compute the states according to eq.˜1, taking $T$ steps. Alternatively, one can cast state evaluation as an optimization problem. While less intuitive, an advantage of this approach is that it admits parallel computation [1, 2, 3]. Depending on the properties of the nonlinear state space model, the optimization algorithm, and the available hardware, the latter approach can be significantly faster than sequential evaluation.
We define the residual and corresponding merit [^2] While minimizing a “merit function” is admittedly counterintuitive, we follow Nocedal and Wright [19, see eq. 11.35] in this convention. function $\mathcal{L}$ by stacking the elements $s_{t}\in\mathbb{R}^{D}$ of a trajectory into a $TD$-dimensional vector $\mathbf{s}$ and considering the vector of temporal differences,
$\mathbf{r}(\mathbf{s})\coloneq\operatorname{vec}\left(\left[s_{1}-f_{1}(s_{0}),\ \ldots,\ s_{T}-f_{T}(s_{T-1})\right]\right)\ \in\ \mathbb{R}^{TD},\qquad\mathcal{L}(\mathbf{s})\coloneq\frac{1}{2}\left\|\mathbf{r}(\mathbf{s})\right\|_{2}^{2},$ (2)
@@
DeepPCR [2] and DEER [1] minimize the merit function using Gauss–Newton updates. Each update takes the form
@@
Due to this block bidiagonal structure, solving $\mathbf{J}(\mathbf{s}^{(i)})^{-1}\,\mathbf{r}(\mathbf{s}^{(i)})$ amounts to solving a linear recursion, which can be done in $\mathcal{O}(\log T)$ time with a parallel scan [5, 20, 21, 1, 3]. Further details are given in Appendix˜A.
@@
Predictability is usually defined through its antonym: unpredictability [9, 10]. In an unpredictable system, the system’s intrinsic sensitivity amplifies small perturbations and leads to massive divergence of trajectories. Predictable systems show the opposite behavior: small perturbations are diminished over time, rather than amplified. The notion of (un)predictability can be formalized through various routes such as chaos theory [22, 23] and contraction analysis [7, 24].
The definition of predictability comes from the Largest Lyapunov Exponent (LLE) [8, 10]:
@@ 3.1 The Merit Function is PL
To begin, we show that the merit function (2) satisfies the Polyak-Łojasiewicz (PL) condition [25], also known as the gradient dominance condition [26]. A function $\mathcal{L}(\mathbf{s})$ is $\mu$-PL if it satisfies, for $\mu>0$,
@@
See Appendix B. This result, known in the literature for general sum-of-squares [27], is included here for context and completeness. ∎
˜1 is important as it characterizes the flatness of the merit function. If $\mu$ is very small in a certain region, this indicates that the norm of the gradient can be very small in that region, which can make gradient-based optimization inefficient. ˜1 also links $\sigma_{\mathrm{min}}(\mathbf{J})$—important for characterizing the conditioning of $\mathbf{J}$—to the geometry of the merit function landscape.
@@ 3.2 Merit Function PL Constant is Controlled by the Largest Lyapunov Exponent of Model
The proof mechanism we have sketched upper and lower bounds $\|\mathbf{J}^{-1}\|_{2}$ in terms of norms of Jacobian products. We only use the assumption in eq.˜10 to express those bounds in terms of $\lambda$. As we discuss at length in Appendix˜C, we can use different assumptions from eq.˜10 to get similar results. ˜2 and its proof should be thought of as a framework, where different assumptions (which may be more or less relevant in different settings) can be plugged in to yield specific results.
@@ 3.3 Residual function Jacobian Inherits the Lipschitzness of the Nonlinear State Space Model
$\forall\,t>1,\ \ \ s,s^{\prime}\in\mathbb{R}^{D}:\quad\|J_{t}(s)-J_{t}(s^{\prime})\|\leq L\|s-s^{\prime}\|,$
@@ 4 Rates of Convergence for Optimizing the Merit Function
Theorem 4 also lets us establish key worst-case and average-case bounds on the number of steps needed for Gauss-Newton to converge to within a given distance of the solution. In particular, when $\chi_{w}$ does not depend on the sequence length $T$, then ˜4 implies Gauss-Newton will only require $\mathcal{O}\left((\log T)^{2}\right)$ total computational time, with one $\log$ factor coming from the parallel scan at each optimization step and the other coming from the total number of optimization steps needed. We elaborate on these points in Appendix G.
@@
See Appendix H. We make no claim about the originality of lower bounding the size of the basin of quadratic convergence in Gauss-Newton. In fact, our proof of ˜5 closely follows the convergence analysis of Newton’s method in Section 9.5.3 of Boyd and Vandenberghe [28]. Our contribution is we highlight the elegant way the predictability $\lambda$ and non-linearity $L$ of a dynamical system influence an important feature of its merit function’s landscape. ∎
@@ 5 Experiments
We conduct experiments to support the theory developed above, demonstrating that predictability enables parallelization of nonlinear SSMs. To illustrate this point, we use Gauss-Newton optimization (aka DEER [1, 2]). We provide more experimental details in Appendix J. Our code is at https://github.com/lindermanlab/predictability_enables_parallelization
@@
Theorem 2 predicts a sharp phase transition in the conditioning of the merit function at $\lambda=0$, which should be reflected in the number of optimization steps required for convergence. To empirically validate this prediction, we vary both the LLE and sequence length $T$ within a parametric family of recurrent neural networks (RNNs), and measure the number of steps DEER takes to converge. We generate mean-field RNNs following Engelken et al. [29], scaling standard normal weight matrices by a single parameter that controls their variance and therefore the expected LLE. In Figure˜2, we observe a striking correspondence between the conditioning of the optimization problem (represented by $\log\tilde{\mu}$, where $\tilde{\mu}$ is the lower bound for $\mu$ from ˜2) and the number of steps DEER takes to convergence. This relationship holds across the range of LLEs, $\lambda$, and sequence lengths, $T$. There is a rapid threshold phenomenon around $\lambda=0$, which divides predictable from unpredictable dynamics, precisely as expected from ˜2. As we discuss in Appendix J.1, the correspondence between $\log\tilde{\mu}$ and the number of optimization steps needed for convergence can be explained by DEER iterates approaching the basin of quadratic convergence with linear rate.
@@
Figure 2: Threshold phenomenon in DEER convergence based on system predictability. In a family of RNNs, DEER has fast convergence for predictable systems and prohibitively slow convergence for chaotic systems. Left (Theory): We depict Theorem 2, illustrating how the conditioning of the optimization problem degrades as $T$ and the LLE ($\lambda$) increase. Center (Experiment): We vary $\lambda$ across the family of RNNs, and observe a striking concordance in the number of DEER optimization steps empirically needed for convergence with our theoretical characterization of the conditioning of the optimization problem. Right: For 20 seeds, each with 50 different values of $\lambda$, we plot the relationship between $\lambda$ and the number of DEER steps needed for convergence for the sequence length $T=1000$ (gray line in left and center panels). We observe a sharp increase in the number of optimization steps at precisely the transition between predictability and unpredictability.
@@
DEER may still converge quickly even if the system is unpredictable in certain regions. As long as the system is predictable on average, as indicated by a negative LLE, DEER can still converge quickly. This phenomenon is why we framed Theorem 2 in terms of the LLE $\lambda$ and burn-in constants $a$, as opposed to a weaker result that assumes the system Jacobians have singular values less than one over the entire state space (see our discussion of condition (10) vs. condition (21) in Appendix˜C).
@@
Notably, prior works such as Lim et al. [1] and Gonzalez et al. [3] initialized optimization from $\mathbf{s}^{(0)}=\mathbf{0}$, which lies entirely in the unstable region. Thus, our theoretical insights into predictability and parallelizability suggest practical improvements for initialization.
Figure 3: DEER converges quickly for Langevin dynamics in a two-well potential. (Left) An illustration of the two-well potential state space in $D=2$. We superimpose a contour plot of the potential on a color scheme showing the spectral norm of the dynamics Jacobian (blue indicates stability, red instability). (Center) A trace plot for the $y$-coordinate. The LLE of the system is 1.45-1.45. (Right) We observe that this system, which has negative LLE, enjoys sublinear scaling in the sequence length $T$ in the number of DEER iterations needed to converge. We plot the median number of DEER steps to convergence over 20 random seeds.
@@
Finally, we demonstrate a practical application of our theory in the efficient parallelization of chaotic observers. Observers are commonly used to reconstruct the full state of a system from partial measurements [30, 31]. On nine chaotic flows from the dysts benchmark dataset [32], Table 1 shows that while DEER converges prohibitively slowly on chaotic systems, it converges rapidly on stable observers of these systems, in accordance with our theory that predictability implies parallelizability. For more details, see Appendix J.5.